# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v2

@@ Appendix I Complete Ideal Functionality for Profile P2
O.1 Mask Properties
Lemma O.1 (Mask Cancellation – Formal Statement).
@@
Lemma O.3 (Mask Hiding – Formal Statement).
@@ Conclusion:
O.2 Irwin-Hall Distribution Analysis
@@
Definition O.5 (Discrete Irwin-Hall Distribution).
@@
Lemma O.6 (Irwin-Hall Moments).
@@
Definition O.9 (Smooth Rényi Divergence [Raccoon2024]).
@@
Lemma O.10 (Security Reduction via Rényi Divergence [BLL+15, Raccoon2024]).
@@
Theorem O.11 (Irwin-Hall Security for Threshold ML-DSA).
@@
$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}(\mathcal{A})\leq R_{2}^{\epsilon}\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}(\mathcal{A})+q_{S}\cdot\epsilon$
where $q_{S}$ is the number of signing queries, and $R_{2}^{\epsilon},\epsilon$ are computed below.
Proof O.12.
@@
4 $8.9\times 10^{-8}$ Negligible
9 $2.3\times 10^{-5}$ Negligible
@@
Conclusion: Over $q_{S}=2^{64}$ signing queries with $|S|\leq 33$:
$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}$ $\leq R_{2}^{\epsilon}\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}+q_{S}\cdot\epsilon$
$\leq 1.05\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}+2^{64}\cdot 10^{-30}$
$\leq 1.05\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}+2^{-36}$
The multiplicative factor ($\leq 1.05$ for $|S|\leq 33$, or $\leq 1.003$ for $|S|\leq 17$) represents at most a 5% security loss, which is negligible for cryptographic purposes. The additive term ($2^{-36}$) is also negligible.
O.3 Lagrange Coefficient Bounds
Lemma O.13 (Lagrange Coefficient Characterization).
@@
Lemma O.14 (Coefficient Magnitude).
@@
Corollary O.16 (Threshold Bound for Short Coefficients).
@@
O.4 Rejection Sampling Correctness
Lemma O.17 (Z-Bound Check Correctness).
@@
Lemma O.18 (R0-Check Correctness).
@@
O.5 Challenge Invertibility
Lemma O.20 (Challenge Invertibility in $R_{q}$).
@@
Corollary O.22 (Security Implication).

### v3

@@ 5 Output
Q.1 Mask Properties
Lemma Q.1 (Mask Cancellation – Formal Statement).
@@
Lemma Q.3 (Mask Hiding – Formal Statement).
@@ Conclusion:
Q.2 Irwin-Hall Distribution Analysis
@@
Definition Q.5 (Discrete Irwin-Hall Distribution).
@@
Lemma Q.6 (Irwin-Hall Moments).
@@
Definition Q.9 (Smooth Rényi Divergence [Raccoon2024]).
@@
Lemma Q.10 (Security Reduction via Rényi Divergence [BLL+15, Raccoon2024]).
@@
Theorem Q.11 (Irwin-Hall Security for Threshold ML-DSA).
@@
$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}(\mathcal{A})\leq\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}(\mathcal{A})+\epsilon}\cdot R_{2}^{\epsilon}+q_{S}\cdot\epsilon$
where $q_{S}$ is the number of signing queries, and $R_{2}^{\epsilon},\epsilon$ are computed below. Since $R_{2}^{\epsilon}\leq 1.05$ for $|S|\leq 33$, this gives $\mathsf{Adv}_{\mathsf{IH}}\leq 1.05\sqrt{\mathsf{Adv}_{\mathsf{U}}}+\mathsf{negl}$.
Proof Q.12.
@@
4 $8.9\times 10^{-6}$ Negligible
9 $2.3\times 10^{-4}$ Negligible
@@
Conclusion: Applying Lemma Q.10 with $\alpha=2$ over $q_{S}=2^{64}$ signing queries with $|S|\leq 33$:
$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}$ $\leq\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}+\epsilon}\cdot R_{2}^{\epsilon}+q_{S}\cdot\epsilon$
$\leq 1.05\cdot\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}+10^{-30}}+2^{64}\cdot 10^{-30}$
$\leq 1.05\cdot\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}}+2^{-36}$
Security interpretation: The bound involves $\sqrt{\mathsf{Adv}_{\mathsf{U}}}$ due to the generic Rényi probability transfer. For ML-DSA-65 with $\mathsf{Adv}_{\mathsf{U}}\leq 2^{-192}$ (NIST Level 3), we have:
$\mathsf{Adv}_{\mathsf{IH}}\leq 1.05\cdot 2^{-96}+2^{-36}<2^{-35}$
This provides approximately 35-bit security margin above any polynomial-time attacker. For practical deployments where $|S|\leq 17$, the tighter bound $R_{2}^{\epsilon}\leq 1.003$ gives security margin $2^{-36}$.
Remark: For Fiat-Shamir with Aborts signatures specifically, tighter reductions exist [Raccoon2024] that avoid the square root loss by exploiting rejection sampling structure. Our conservative bound suffices for the claimed security level.
Q.3 Lagrange Coefficient Bounds
Lemma Q.13 (Lagrange Coefficient Characterization).
@@
Lemma Q.14 (Coefficient Magnitude).
@@
Corollary Q.16 (Threshold Bound for Short Coefficients).
@@
Q.4 Rejection Sampling Correctness
Lemma Q.17 (Z-Bound Check Correctness).
@@
Lemma Q.18 (R0-Check Correctness).
@@
Q.5 Challenge Invertibility
Lemma Q.20 (Challenge Invertibility in $R_{q}$).
@@
Corollary Q.22 (Security Implication).